What will happen in the time recorded by a pendulum clock when temperature increases and decreases?

Grade:11

1 Answers

jagan
6 years ago
when a pendulum clock is heated up it’s length increses so it will take more time to complete the oscillation hence it looses the time on heating. it gains the time on cooling
       \Delta T= \frac{1}{2}\alpha \Delta \theta T